# Single Image Test-Time Adaptation for Segmentation

### v1

@@ 3 Methods
The following TTA methods are considered in our experiments: Entropy-Minimization (Ent) is a method based on entropy minimization inspired by semi-supervised learning proposed by [43]. Pseudo-Labelling (PL) is training with pseudo-labels obtained from the pretrained segmentation model. Ref is training with pseudo-labels obtained by a dedicated refinement module that takes masks as inputs and outputs a refined mask. dIoU is similar to Ref. However, a single scalar quality estimate is learnt and minimized at test-time. AugCo is the method of [29] based on training for consistency between the prediction and prediction on augmented views adapted to the single-image scenario. Adv is the method proposed by [27], adapted to single image segmentation setup.
The rest of this section first describes TTA with mask refinement in more detail, including the novel adversarial training of the refinement module and then the proposed evaluation metrics. The other methods evaluated in the paper are simple adaptations of existing baselines to the single-image setup without assuming the existence of batch-normalization layers. A description of these baselines, as well as the details of the necessary modifications, is in Appendix B.
@@
TTA with mask refinement is based on the idea that since the output space changes much less than the input space, a mask translation module can be learnt to refine mask predictions on images from target distribution to resemble the source distribution masks. This is similar to domain adaptation methods based on learning a discriminator between source and target domain, with two key differences: Instead of a binary discriminator, a mask refinement module is learnt, so the output is not a scaler but a new segmentation mask. Also, the target domain is not known in advance and can not be used to train the refinement network. At test-time, the refinement network can be viewed as an enhanced pseudo-label generation method.
@@
To train the refinement network $f_{\text{R}}$, we assume access to images from the source distribution and the pretrained segmentation network $f^{\theta}_{\text{S}}(x)$. Given an image $x$ and $x^{\prime}$ generated from $x$ by synthesizing domain shift without changing the label, let us denote as $s=f^{\theta}_{\text{S}}(x)$ and $s^{\prime}=f^{\theta}_{\text{S}}(x^{\prime})$ the corresponding segmentation masks. Then, $f_{\text{R}}$ is trained to predict $m$, given $m^{\prime}$ as input. Predicted masks $m$ can be replaced with ground truth in training time.
An overview of the training pipeline is in Figure 1.
Figure 1: Mask refiner training. The segmenter outputs masks from a training image and a corrupted image simulating domain shift. The mask refiner is then trained to predict the clean image mask given the corrupted image mask as input only - no gradients flow to the segmenter. Visualized on single class prediction.
@@
$\theta_{i}^{\star}=\operatorname*{arg\,min}_{\theta_{i}}\mathcal{L}_{\text{IoU}}(f_{\text{R}}(s_{i}),s_{i})$ (1)
@@
where $\theta_{i}$ are the learnable parameters of $f^{\theta}_{\text{S}}(x)$ and $s_{i}=f^{\theta_{i}}_{\text{S}}(x)$ at iteration $i$.
@@
Refinement module training requires generating masks resembling those that the model would output under domain shift. Instead of swapping input image patches, a heuristic method with many hyper-parameters that could lead to unrealistic artifacts in the masks, this work simulates the mask corruption by using mask predictions on the images from the first few iterations of the Projected Gradient Descent (PGD) [24, 17] adversarial attack, using the inverted mask as target. The more iterations of the attack, the higher the mask corruption. This approach is illustrated in Appendix C. In the first iterations, the most challenging pixels for the network are converted. Similarly, those image areas could be easily impacted by domain shift.
@@
Evaluation metrics The standard semantic segmentation evaluation metric is the mean Intersection over Union (mIoU), where the IoU score of each class is computed from predictions aggregated over the whole dataset
$\text{mIoU}=\frac{1}{\text{C}}\sum_{c=1}^{\text{C}}\text{IoU}_{c}(m_{c},g_{c})$ (2)
@@
where $m_{c},g_{c}$ are we predictions and ground truth values for class $c$ for all pixels across all images and $|\text{I}_{c}|$ is the total number of pixels. This metric does not consider the size of objects or the difficulty of individual images. Finally, per-image results can not be compared.
@@
Two additional metrics are introduced to account for the limitations of the standard mIoU. The first metric is designed to consider class imbalance and difficulty of individual images, focusing on per-class performance. It will be referred to as $\text{m}\overline{\text{IoU}}_{c}$ and can be computed as
$\text{m}\overline{\text{IoU}}_{c}=\frac{1}{\text{C}}\sum_{c=1}^{\text{C}}\frac{1}{|\text{I}_{c}|}\sum_{i\in\text{I}_{c}}\text{IoU}(m_{ic},g_{ic})$ (3)
@@
where $\text{I}_{c}$ is the set of images in which either the prediction or the ground truth mask contains class $c$.
@@
The second metric is focused on per-image performance and can be computed for a single image. It will be referred to as $\text{m}\overline{\text{IoU}}_{i}$ and can be computed as
$\text{m}\overline{\text{IoU}}_{i}=\frac{1}{|\text{I}|}\sum_{i\in\text{I}}\frac{1}{|\text{C}_{i}|}\sum_{i\in\text{C}_{i}}\text{IoU}(m_{ic},g_{ic})$ (4)
where $\text{C}_{i}$ is the set of classes in the predicted masks and the ground truth. Since this work focuses on single-image adaptation and the analyses often require per-image performance, this is the metric reported in most experiments. A similar metric was employed by [41] where the score is determined by averaging the score for the classes in the ground truth only, leading to more optimistic results.
@@
Figure 2: GTA5-C error reduction (%) depending on corruption levels. TTA with overall optimal hyper-parameters for COCO-C.
@@
Figure 3: The relationship between per-image scores before and after adaptation on the GTA5-C dataset. The difference between non-adapted (NA) $\text{m}\overline{\text{IoU}}_{i}$ and the $\text{m}\overline{\text{IoU}}_{i}$ after TTA is shown ($\text{m}\overline{\text{IoU}}_{i}\Delta$). A least-squares line fitted to the points in yellow.
Figure 4: The relationship between per-image non-adapted (NA) prediction entropy and TTA $\text{m}\overline{\text{IoU}}_{i}$ improvement on the GTA5-C dataset. The difference between NA $\text{m}\overline{\text{IoU}}_{i}$ and the TTA $\text{m}\overline{\text{IoU}}_{i}$ is shown ($\text{m}\overline{\text{IoU}}_{i}\Delta$). Least-squares line fitted to the points in yellow.
@@
4 Experiments
The TTA methods are evaluated on two semantic segmentation models pretrained on the GTA5 and COCO datasets. Experiments with each model consist of two parts: 1) Hyper-parameter tuning and analysis on a synthetic validation set and 2) evaluation on real-world datasets with domain shift.
The validation set is based on a set of 40 images from the segmentation model training dataset extended with a diverse set of 9 synthetic corruptions at three corruption levels from [12] such as blur, noise or fog, simulating different domain shifts. Details about the corruptions can be found in Appendix D. These extended datasets based on the GTA5 and COCO datasets are referred to as GTA5-C and COCO-C, respectively. Each validation dataset consists of 1200 images in total. For each TTA method, optimizing either all the network parameters or normalization parameters only is considered, resulting in at least two different setups for each method. Further, when applicable, the cross-entropy and IoU losses are compared. This results in four setups for the Ref, PL and AugCo methods.
Shared implementation details The refinement network architecture is a U-Net [32] with an EfficentNet-B0 [38] backbone pre-trained on ImageNet from the Timm library [46]. It is trained with the AdamW [23] optimizer with a learning rate of $1e^{-3}$ and the Cross-Entropy (CE). The SGD optimizer is used for TTA and the hyper-parameters considered during validation are the learning rate and number of TTA iterations. The maximum possible number of iterations is 10 to limit the computational requirements. Early experiments with AdamW for TTA showed high divergence rate.
@@ 4 Experiments
Figure 5: COCO-C error reduction (%) depending on corruption levels. TTA with overall optimal hyper-parameters for COCO-C.
4.1 GTA5 $\rightarrow$ Cityscapes, ACDC
The GTA5-pretrained model is the best-performing model of [41] (DeepLabV2). The results of different methods with parameters selected for the overall best performance across all corruptions and levels can be found in Table 2. It can be observed that largely improved results are achieved either by PL with IoU loss, optimizing normalization parameters only, or by Ref with IoU loss, optimizing all the parameters. The best-performing method is Ref, improving by 3.51 % over the non-adapted baseline (NA), reducing the total segmentation error by 5.41 %. Other methods only marginally improve over NA or show no improvement at all. Optimizing CE generally yields worse results than IoU. While updating normalization parameters may only stabilize Ent, optimizing all the parameters for optimal performance of Ref is important. For other methods, the difference is small, yet optimizing normalization parameters only is faster and thus recommended.
In Figure 2, the total error reduction results with a single set of optimal hyper-parameters for each method are reported for each corruption level and kind. Results show that it is not possible to find a single set of hyper-parameters that would perform well across all the corruption levels with current methods. While all methods improve performance on level 1 corruptions, from level 3, negative results can be observed for some of the corruptions, and all methods yield negative results on level 5. Ref outperforms the other methods on the majority of corruption kinds and corruption levels. Aggregated results across all corruptions show that the negative results for level 5 and mixed results for level 3 are mostly outweighed by the gains on level 1, resulting in overall positive results. This analysis suggests that unless it is known in advance what kind of corruption or corruption levels will be present after deployment, strategies on method and hyper-parameter selection for each image should be considered.
In Appendix E, it is shown that if one could select optimal hyper-parameters for each method, corruption kind and level, results would improve substantially. Moreover, Ref significantly outperforms the other methods on most corruptions, the blur corruption kinds being a notable exception. Significant improvements can be observed compared to other methods, especially on different kinds of noise and jpeg corruption at level 5.
Only the methods with positive TTA results are used for further analysis, namely Ref, PL, AugCo and Ent. Finally, the relationship between the non-adapted (NA) performance and the performance improvement on individual images for different methods is visualized in Figure 3. The analysis shows Ref outperforms other methods, especially on images that had low initial $\text{m}\overline{\text{IoU}}_{i}$, while the performance of PL is consistent across all initial scores yet not as powerful for initial low scores. While Ent makes performance worse for low initial scores and improves more as the initial score increases, AugCo shows consistent improvements across all initial scores similarly to PL, yet to a smaller extent.
Figure 6: TTA $\text{m}\overline{\text{IoU}}_{i}$ error reduction (%) of GTA5 model on the driving test datasets with hyper-parameters selected on GTA5-C.
If the $\text{m}\overline{\text{IoU}}_{i}$ for each image were known, it could be used to either select a method performing best on those values or to select hyper-parameters. In Figure 4, an analogous analysis is performed, replacing the $\text{m}\overline{\text{IoU}}_{i}$ with segmentation prediction entropy. Similar results as with the $\text{m}\overline{\text{IoU}}_{i}$ can be observed, showing that the segmentation prediction entropy is a good proxy for initial segmentation $\text{m}\overline{\text{IoU}}_{i}$.
After the best hyper-parameters for each method are selected on the validation set, the methods are evaluated on 5 test datasets: ACDC-Rain, ACDC-Fog, ACDC-Night and ACDC-Snow, and Cityscapes. The Cityscapes represents a domain shift from synthetic to real images. The first four datasets are created by splitting the ACDC dataset by different conditions and, apart from the synthetic to real, further represent adverse weather conditions domain shift. Each test set consists of 500 images.
The test results are reported in Figure 6. Similar to validation, the Ref and PL methods perform best across all datasets. While not outperforming Ref on all the datasets, the performance of PL is consistently better than the other methods while Ref is outperformed or matched by the other methods on Cityscapes and ACDC-night.
4.2 COCO $\rightarrow$ VOC
The segmentation model is an official Torchvision DeepLabV3 with Resnet50 backbone trained on the COCO dataset with a subset of 20 VOC classes. The results of different methods with parameters selected for the overall best performance across all corruptions and levels can be found in Table 3. Major improvements are obtained by the PL and Ref methods. PL outperforms Ref, in contrast to GTA5-C experiments. The best improvement is by 3.28 %, reducing the total segmentation error by 7.3 %. Again, best results for each method are always achieved by IoU, outperforming CE in all cases. In contrast to GTA5-C, Ent achieves better results when optimizing all the network parameters. The same holds for PL. For Ref, optimizing all the parameters is important. For most methods, the difference is small; overall, improvements over the non-adapted baseline are marginal.
@@ 4.2 COCO $\rightarrow$ VOC
Figure 7: The relationship between per-image scores before and after adaptation on the COCO-C dataset. The difference between non-adapted (NA) $\text{m}\overline{\text{IoU}}_{i}$ and the $\text{m}\overline{\text{IoU}}_{i}$ after TTA is shown ($\text{m}\overline{\text{IoU}}_{i}\Delta$). A least-squares line fitted to the points in yellow.
Figure 8: The relationship between per-image non-adapted (NA) prediction entropy and TTA $\text{m}\overline{\text{IoU}}_{i}$ improvement on the COCO-C dataset. The difference between NA $\text{m}\overline{\text{IoU}}_{i}$ and the TTA $\text{m}\overline{\text{IoU}}_{i}$ is shown ($\text{m}\overline{\text{IoU}}_{i}\Delta$). Least-squares line fitted to the points in yellow.
@@
The total error reduction results with a single set of optimal hyper-parameters for each method are reported for each corruption level and kind in Figure 5. The results slightly differ from those for the GTA5-C, as in this case, PL outperforms Ref. Both methods are again consistently better than the other methods, but positive results are reported for most corruptions on both level 1 and level 3.

### v2

@@ 4 Methods
In total, six different methods are implemented and evaluated, including traditional TTA baselines, methods from other tasks and novel methods. All the methods consist of optimizing a self-supervised loss, the specifics of the loss being what differentiates the methods. It can be formalized as follows:
$\theta_{i+1}=\arg\min_{\theta_{i}}\mathcal{L}(f_{S}^{\theta_{i}},x)$
where $x$ is the input image, $\theta_{i}$ are the parameters of the segmentation network $f_{S}^{\theta_{i}}$at the $i$-th iteration and $\mathcal{L}$ is the self-supervised loss function.
The methods considered are:
  • Entropy-Minimization (Ent), a method proposed by Wang et al. (2020b) inspired by semi-supervised learning where the self-supervised objective is the prediction entropy. It has been used as a baseline by most of the TTA work. Only normalization layer parameters are typically updated to reduce the computational cost. Whether batch normalization statistics are updated or not varies.
  • Pseudo-Labelling (PL), also commonly referred to as self-training. The model is finetuned with pseudo-labels obtained from the pretrained segmentation model. There are many improvements and modifications. The standard approach is to threshold the predicted probabilities and only train the model on the most confident predictions.
  • Augmentation-Consistency (AugCo), proposed by Prabhu et al. (2021), is a method based on self-training enhanced by also optimizing for consistency between the original prediction and the prediction on augmented views, adapted to the single isolated image scenario.
  • Adversarial-Attack (Adv) is the method proposed by Nguyen et al. (b) for image classification TTA, adapted to the single, isolated image segmentation.
  • Mask Refinement (Ref) is one of the proposed methods and can be considered an enhanced pseudolabelling method. The pseudo-labels are obtained by a learnt refinement module that takes logit masks as inputs and outputs a refined segmentation mask. A method based on this idea has been implemented in medical imaging Karani et al. (2021) but never tested on non-medical tasks.
  • Deep-Intersection-over-Union (dIoU), the second proposed method is similar to Ref. However, a single-scalar quality estimate is predicted by a learnt module and minimized at test time. It is similar to using a GAN-like discriminator, as done in unsupervised domain adaptation literature Tzeng et al. (2017).
Only the necessary modifications to make the methods applicable to the single, isolated image segmentation setup with no assumptions of specific network architecture were applied to the existing methods.
The proposed methods based on learnt mask-refinement and mask-discriminator modules will be described in the rest of this section. A description of other methods and the details of their modifications in this work are in Appendix B.
The method proposed in this work is presented first and the differences from previous work are detailed aftwerwards.
@@
TTA with mask refinement is based on the idea that since the output space changes much less than the input space, a mask translation module can be learnt to refine mask predictions on images from target distribution to resemble the masks obtained from source images. At test time, the refinement network can be viewed as an enhanced pseudo-label generation method. These pseudo-labels can then be used both as supervision for the segmenter or a direct replacement of the segmentation output without any parameter optimization. However, the second option is unlikely to tackle highly distorted masks since the refined mask cannot improve gradually.
@@
To train the refinement network $f_{\text{R}}^{\phi}$ with learnable parameters $\phi$, images from the source distribution and the pretrained segmentation network $f^{\theta}_{\text{S}}$ are required. Given an image $x$ and $x^{\prime}$ generated from $x$ by synthesizing a covariate domain shift (not changing the label), let us denote as $s=f^{\theta}_{\text{S}}(x)$ and $s^{\prime}=f^{\theta}_{\text{S}}(x^{\prime})$ the corresponding segmentation masks. Then, $f_{\text{R}}$ is trained to predict $s$, given $s^{\prime}$ as input:
$\operatorname*{arg\,min}_{\phi}\mathcal{L}_{\text{CE}}(f_{\text{R}}^{\phi}(s^{\prime}),s)$ (1)
Predicted masks $s$ can also be replaced with ground truth $g$ at training time:
$\operatorname*{arg\,min}_{\phi}\mathcal{L}_{\text{CE}}(f_{\text{R}}^{\phi}(s^{\prime}),g)$ (2)
where $\mathcal{L}_{\text{CE}}$ is the cross-entropy loss.
@@
$\theta_{i+1}=\operatorname*{arg\,min}_{\theta_{i}}\mathcal{L}_{\text{IoU}}(f_{\text{R}}(\overline{f}^{\theta_{i}}_{\text{S}}(x)),f^{\theta_{i}}_{\text{S}}(x))$ (3)
@@
where $\theta_{i}$ are the learnable parameters of $f^{\theta}_{\text{S}}$ at optimization iteration $i$ and $\overline{f}_{S}$ denotes no gradient flow throughout the computations of $f_{\text{S}}$.
An overview of the training pipeline and the TTA with mask-refinement is in Figure 2.
Figure 2: Mask refiner training (left) and Mask Refinement (Ref) TTA inference (right). During training, the segmenter outputs masks from a training image and a corrupted version of the training image simulating domain shift. The mask refiner is then trained to predict the clean image mask given the corrupted image mask as input only - no gradients flow back to the segmenter. At inference time, the segmenter output is fed into the refiner model. The refined output is then used as a pseudo-label to finetune the segmenter. A single gradient update is performed in each TTA iteration, then the masks are updated. The segmenter output may change with the updated weights, which in turn results in a new, possibly better, pseudo-label from the refiner. Visualized on single class prediction.
@@
Refinement module training requires generating masks resembling those that the model would output under domain shift. Since TTA assumes the domain shift is not known in advance, the goal is to synthetically generate diverse realistic segmentation masks representing masks predicted on images with domain shift. The advantage of the refinement module is that only the output space corrupted masks are needed. It doesn’t matter how these were obtained since the refinement module is independent of the input images. This work simulates the mask corruptions by using mask predictions on the images from the first few iterations of a Projected Gradient Descent (PGD) Madry et al. (2017); Kurakin et al. (2018) adversarial attack, using the inverted mask as target. The more iterations of the attack, the higher the mask corruption, but the less realistic it becomes. Examples of generated corrupted masks are shown in Appendix C. The intuition behind this adversarial approach is that in the first iterations, the most challenging pixels for the network are converted. Similarly, those image areas could be easily impacted by domain shift. The idea of using adversarial attack as a representative of domain shifts was also used in Nguyen et al. (b) and it was previously shown that adversarial robustness improves robustness to some domain shift kinds Croce et al. (2020).
TTA with Deep-Intersection-over-Union is almos equivalent to Mask Refinement (Ref). The only difference is that rather than learning a refined mask, the network is optimized to predict the value of the IoU loss. The IoU loss can be computed from ground-truth at training time but the network receives the predicted mask as input only.
Difference from unsupervised domain adaptation literature: There are many domain adaptation methods based on learning a discriminator between source and target domain. The key difference is that in this work, a module estimating how corrupted the mask is or a mask refinement module is learnt, instead of a binary classifier. Hence, the output is either a new segmentation mask, or a real number, rather than a class (domain) probability. This has the advantage that the mask needs to be considered as a whole, rather than allowing the network to focus on the most discriminative parts only. Also, the target domain is not known in advance and cannot be used to train the refinement network - masks resembling the target domain masks need to be synthesized.
Difference from medical image segmentation TTA with denoising autoencoders The idea of learning to predict a refined mask in the output space is the same as in in Karani et al. (2021). The main differences stem from the domain difference, since their approach is tailored to the medical domain. In Karani et al. (2021), the corrupted masks are obtained through swapping input image patches, a heuristic method with many hyper-parameters. Such augmentation makes sense when texture is more important than shapes, which is often not the case outside the medical domain. This work proposes exploiting adversarially-attacked images to synthesize corrupted masks. Further, this work excludes methods that make changes to the segmentation model architecture while Karani et al. (2021) rely on an additional normalization module. Their module architecture is designed with domain shifts such as intensity transformations, typical for the medical domain, in mind. Updates of parameters in the segmentation network itself are not allowed during TTA.
5 Experiments
The structure of this section is as follows:
  1. 1.
    Evaluation metrics: Given the focus of this study on SITTA and per-image performance analysis, we underscore the need for an image-level evaluation metric. The widely used mean Intersection over Union (mIoU) metric is typically applied at the dataset level and its adaptation for image-level assessment is not standardized.
  2. 2.
    Experimental Setup: Experiment settings shared across experiments such as network architectures or hyper-parameters. Creation of the synthetic SITTA training set derived from the segmentation training dataset is also explained.
  3. 3.
    Experimental Results and Analysis: Experiment results and analysis. The TTA methods are evaluated on two semantic segmentation models pretrained on the GTA5 Richter et al. (2016) and COCO Lin et al. (2014) datasets.
5.1 Evaluation metrics
@@
The standard semantic segmentation evaluation metric is the mIoU, where the IoU score of each class is computed from predictions aggregated over the whole dataset and the per-class scores are averaged
$\text{mIoU}=\frac{1}{\text{C}}\sum_{k=1}^{\text{C}}\text{IoU}_{k}(m_{k},g_{k})$ (4)
@@
where $m_{k},g_{k}$ are the predictions and ground truth values for class $k$ for all pixels across all images. Concatenating all the masks into a single one and then computing the metric would not change the results, each pixel has the same weight independent of the image size or difficulty. This metric does not consider the size of objects or the difficulty of individual images. Per-image results cannot be compared, since not all classes are typically present in an image an it is not clear what value the score for that class should be.
@@
Two additional metrics are introduced to account for the limitations of the standard mIoU and make the evaluation more fine-grained. The first metric is designed to consider class imbalance and difficulty of individual images, focusing on per-class performance. It will be referred to as $\text{m}\overline{\text{IoU}}_{c}$ and is defined as
$\text{m}\overline{\text{IoU}}_{c}=\frac{1}{\text{C}}\sum_{k=1}^{\text{C}}\frac{1}{|\text{I}_{k}|}\sum_{i\in\text{I}_{k}}\text{IoU}(m_{ik},g_{ik})$ (5)
@@
where $\text{I}_{k}$ is the set of images in which either the prediction or the ground truth mask contains class $k$ and $|\text{I}_{k}|$ is the total number of images in $\text{I}_{k}$.
@@
The second metric is focused more on per-image performance and can be computed for a single image. It will be referred to as $\text{m}\overline{\text{IoU}}_{i}$ and is defined as
$\text{m}\overline{\text{IoU}}_{i}=\frac{1}{|\text{I}|}\sum_{n\in\text{I}}\frac{1}{|\text{C}_{n}|}\sum_{k\in\text{C}_{n}}\text{IoU}(m_{nk},g_{nk})$ (6)
where $\text{C}_{n}$ is the set of classes in the predicted masks or the ground truth of image $n$. $I$ denotes the set of all images. This is the metric reported in our experiments unless stated otherwise. It allows for per-image performance comparison with the disadvantage of not accounting for class imbalance - less frequent classes (on the image level) get smaller weight.
Similar metrics were recently considered by other works Volpi et al. (2022), typically only aggregating over images where the given class appears in the ground truth (as opposed to either the ground truth or the prediction). This has the advantage that mistakes are only accounted for once, making the metric more optimistic than ours. On the other hand, information about the errors is lost since the error is only computed for the ground truth class independently of what the incorrectly predicted class is.
On test datasets, the mDice (mean over the per-class dice scores Milletari et al. (2016)) and Accuracy (overall percentage of correctly classified pixels, regardless of class) metrics are also reported.
5.2 Experiment setting
SITTA training set. The SITTA training set for each model is derived from a set of 40 images from the segmentation model’s training dataset extended with a set of 9 synthetic corruptions at three severity levels from Hendrycks & Dietterich (2019) such as blur, noise or fog, simulating different domain shifts. The original images are also included since the TTA methods should not harm the model on source domain images. Details about the corruption can be found in Appendix D. These synthetic datasets based on the GTA5 and COCO datasets are referred to as GTA5-C and COCO-C, respectively. Since the original images without corruption are also included, each SITTA training dataset consists of 1200 images (40 images, 9 + 1 corruption, three corruption levels).
SITTA hyper-parameters. For each TTA method, optimizing all the network parameters or normalization parameters is only considered, resulting in at least two different setups for each method. Further, when applicable (the methods compute a segmentation loss based on masks, as opposed to another self-supervised loss such as the prediction entropy), the CE and IoU losses are compared. While training segmentation models with a loss that takes class imbalance into account, such as the CE and the Dice loss Milletari et al. (2016), is standard, TTA work on image segmentation has relied on cross-entropy, which is also suboptimal from the point of view of the evaluation metric. It is desirable to align the optimization metric with the evaluation metric as much as possible. This results in four setups for the Ref, PL, and AugCo methods. The learning rate and number of TTA iterations are considered from learning hyper-parameters. The maximum possible number of iterations is 10 to limit the computational requirements. Reasonable learning rate values are found via a grid search and then extended with other promising values based on the initial results.
Shared implementation details. The refinement network architecture is a U-Net Ronneberger et al. (2015) with an EfficentNet-B0 Tan & Le (2019) backbone pre-trained on ImageNet from the Timm library Wightman (2019). It is trained with the AdamW Loshchilov & Hutter (2017) optimizer with a learning rate of $1e^{-3}$ and the Cross-Entropy (CE) loss. The SGD optimizer is used for the TTA since early experiments with AdamW showed a high divergence rate.
5.3 Experiment results
@@
Figure 3: GTA5-C $\text{m}\overline{\text{IoU}}_{i}$ error reduction (%) depending on corruption levels. TTA with overall optimal hyper-parameters for GTA5-C.
(a) $\text{m}\overline{\text{IoU}}_{i}$ NA
(b) entropy NA
@@
Figure 4: The relationship between per-image scores (a) or entropy (b) before and the score after adaptation on the GTA5-C dataset. The difference between non-adapted (NA) $\text{m}\overline{\text{IoU}}_{i}$ or entropy and the $\text{m}\overline{\text{IoU}}_{i}$ after TTA is shown ($\text{m}\overline{\text{IoU}}_{i}\Delta$). A least-squares line fitted to the points is shown in yellow.
GTA5 $\rightarrow$ Cityscapes, ACDC. This experiment explores the performance of the TTA methods on a model trained on a synthetic driving dataset, GTA5, evaluating on real-world driving datasets under different weather conditions. The GTA5-pretrained model is the best-performing model of Volpi et al. (2022) (DeepLabV2).
Since current methods do not consider different hyper-parameters for individual images, a single set with overall best performance across all corruptions and corruption levels is selected. The aggregated results with these overall optimal hyper-parameters on the SITTA training set can be found in Table 1. It can be observed that the biggest improvements are achieved either by PL with IoU loss, optimizing normalization parameters only, or by Ref with IoU loss, optimizing all the parameters. The best-performing method is Ref, improving by 3.51 % over the non-adapted baseline (NA). Other methods only marginally improve over NA or show no improvement at all. Optimizing CE generally yields worse results than optimizing the IoU. While updating normalization parameters only may stabilize Ent, optimizing all the parameters is essential for optimal performance of Ref. For other methods, the difference is smaller - optimizing normalization parameters only is faster and thus recommended.
In Figure 3, the total error reduction results with the same set of overall optimal hyper-parameters for each method are shown but for each corruption level and kind separately. It can be seen that it is not possible to find a single set of hyper-parameters that would perform well across all the corruption levels with these methods. While all methods except for AugCo improve performance on level 1 corruptions, from level 3, negative results can be observed more often and many methods deteriorate/only slightly improve on level 5. Ref outperforms the other methods on the majority of corruption kinds and corruption levels. The aggregated results across all corruptions showed that the negative results are outweighed by the gains on level 1, resultin g in overall positive results for most of the methods.
In Figure 5, it is shown that if one could select optimal hyper-parameters for each corruption kind and level, results would improve substantially for many of the corruptions. This analysis suggests that unless the domain shift is known in advance, strategies with method and hyper-parameter selection for each image separately should be explored.
Figure 5: GTA5-C error reduction difference (%) between overall optimal hyperparameters and hyper-parameters selected for each corruption kind separately. The hyper-parameters were selected on GTA5-C.
Only the methods with overall positive TTA results are considered for further analysis, namely Ref, PL, AugCo, and Ent. The relationship between the non-adapted (NA) performance and the performance improvement on individual images for different methods is visualized in Subfigure 4(a). The analysis shows Ref outperforms other methods, especially on images that had low initial $\text{m}\overline{\text{IoU}}_{i}$, while the performance of PL is more consistent across all initial scores but not as powerful for initial low scores. While Ent makes performance worse for low initial scores and improves more as the initial score increases, AugCo shows consistent improvements across all initial scores similarly to PL but to a smaller extent.
(a) GTA5 model
(b) COCO model
Figure 6: TTA $\text{m}\overline{\text{IoU}}_{i}$ error reduction (%) on test datasets with hyper-parameters select on the TTA training datasets.
If the $\text{m}\overline{\text{IoU}}_{i}$ for each image were known (the ground truth is necessary for its computation), it could be used to either select a method performing best on those values or to select hyper-parameters. In Subfigure 4(b), an analogous analysis is performed, replacing the $\text{m}\overline{\text{IoU}}_{i}$ with segmentation prediction entropy, which does not require any supervision. Similar results as with the $\text{m}\overline{\text{IoU}}_{i}$ can be observed.
After selecting the best hyper-parameters for each method on the SITTA training set, the methods are evaluated on 5 test datasets: ACDC-Rain, ACDC-Fog, ACDC-Night, ACDC-Snow, and Cityscapes. The Cityscapes represent a domain shift from synthetic to real images, while ACDC datasets add adverse weather conditions, making the domain shift even greater. The first four datasets are created by splitting the ACDC dataset by different conditions. Each of the test sets consists of 500 images. The test results are reported in Figure 6(a) and Table 2. Similarly to SITTA training datasets, either Ref or PL methods perform best, depending on the dataset. While not outperforming Ref on all the datasets, the performance of PL is consistently better than the other methods while Ref is outperformed or matched by the other methods on Cityscapes and ACDC-night. Table 2 shows that in standard metrics that do not consider per-image difficulty, AugCo outperforms the other methods in many cases.
method
dataset metric NA Ref PL Ent AugCo
cityscapes $m\overline{\mathrm{IoU}}_{i}$ 34.40 34.72 37.14 35.12 35.52
$m\overline{\mathrm{IoU}}_{c}$ 28.71 28.65 30.70 29.09 29.53
mIoU 42.90 40.37 43.28 43.41 44.05
mDice 55.78 52.53 56.04 56.17 56.89
Accuracy 86.54 83.92 87.15 87.30 87.50
acdc-fog $m\overline{\mathrm{IoU}}_{i}$ 32.03 35.99 35.68 33.92 31.82
$m\overline{\mathrm{IoU}}_{c}$ 24.87 27.29 27.52 26.00 24.69
mIoU 37.65 38.36 39.39 39.15 37.61
mDice 51.42 51.12 53.29 52.68 51.51
Accuracy 73.23 84.11 75.51 75.91 66.63
acdc-night $m\overline{\mathrm{IoU}}_{i}$ 13.60 14.12 15.09 14.13 14.15
$m\overline{\mathrm{IoU}}_{c}$ 10.77 10.96 11.53 10.68 11.01
mIoU 15.79 13.77 15.11 15.53 16.25
mDice 24.38 21.29 23.49 23.74 24.98
Accuracy 52.09 52.47 51.86 52.82 53.09
acdc-rain $m\overline{\mathrm{IoU}}_{i}$ 33.52 35.60 37.16 35.05 34.50
$m\overline{\mathrm{IoU}}_{c}$ 26.15 27.40 28.47 26.89 26.73
mIoU 36.93 36.21 37.61 37.53 37.98
mDice 48.92 47.52 49.44 49.23 50.25
Accuracy 84.56 84.22 85.65 86.13 84.74
acdc-snow $m\overline{\mathrm{IoU}}_{i}$ 31.54 35.60 34.15 31.87 31.81
$m\overline{\mathrm{IoU}}_{c}$ 25.28 28.09 27.16 25.38 25.45
mIoU 35.30 37.89 36.64 35.34 35.40
mDice 48.46 50.35 50.00 48.51 48.53
Accuracy 73.17 81.52 74.48 73.66 73.45
The inconsistencies of results between SITTA training and test suggest that unless the domain shift conditions are known in advance, it is difficult to select hyper-parameters based on a general SITTA training set.
COCO $\rightarrow$ VOC. In this experiment, the performance of TTA methods is studied on a model trained on the COCO dataset and evaluated on the VOC dataset. The segmentation model is an official Torchvision DeepLabV3 model with a Resnet50 backbone trained on the COCO dataset with a subset of 20 VOC classes. In contrast to previous experiments, it is a real-to-real dataset domain shift. The results of different methods with parameters selected for the overall best performance across all corruptions and levels can be found in Table 3. Biggest improvements are obtained by the PL and Ref methods. PL outperforms Ref, in contrast to the GTA5-C experiments. The best improvement is by 3.28 %, reducing the total segmentation error by 7.3 %. Consistently with previous experiments, best results are achieved with the IoU loss, outperforming CE in all cases. In contrast to GTA5-C, Ent achieves better results when optimizing all the network parameters, as opposed to optimizing only the normalization layer parameters. The same holds for PL. For Ref, optimizing all the parameters is again important. Other methods’ improvements over the non-adapted baseline are marginal.
@@
Figure 7: COCO-C $\text{m}\overline{\text{IoU}}_{i}$ error reduction (%) depending on corruption levels. TTA with overall optimal hyper-parameters for COCO-C.
(a) $\text{m}\overline{\text{IoU}}_{i}$ NA
(b) entropy NA
@@
Figure 8: The relationship between per-image scores (a) or entropy (b) before and the score after adaptation on the COCO-C dataset. The difference between non-adapted (NA) $\text{m}\overline{\text{IoU}}_{i}$ or entropy and the $\text{m}\overline{\text{IoU}}_{i}$ after TTA is shown ($\text{m}\overline{\text{IoU}}_{i}\Delta$). A least-squares line fitted to the points is shown in yellow.
@@
The total error reduction results with a single set of optimal hyper-parameters for each method are reported for each corruption level and kind in Figure 7. The results slightly differ from those for the GTA5-C, as in this case, PL is consistently the best method, only rarely outperformed by Ref or AugCo. Negative results on some corruption kinds are already reported for level 1 corruptions.